What is the probability of both coins landing on heads?


1
Expert's answer
2022-06-10T02:32:12-0400

Solution:

There are two coins, both have one head and one tail. So, the probability of one coin on head is 1/2. Two evens must be happened in the same time for our case. For that we need to multiply their probabilities landing on head.

"P(AB)=P(A)P(B)=\frac{1}{2}\times\frac{1}{2}=\frac{1}{4}."

Answer:

"P(AB)=\frac{1}{4}."